OpenStudy (anonymous):

Select the best possible first step to solving the following system by first eliminating the x variable. 2x + 5y = 1 x + 6y = –10

OpenStudy (anonymous):

Multiply the second equation by 2. Multiply the first equation by 2. Multiply the first equation by –2. None of the above

OpenStudy (2bornot2b):

2x + 5y = 1 -2x - 12y = 20

OpenStudy (2bornot2b):

-8y=21

OpenStudy (2bornot2b):

i.e. Multiply the first equation by –2.

OpenStudy (mertsj):

Multiply the second equation by -2

OpenStudy (anonymous):

would the answer be none of the above? because none of those are choices

OpenStudy (anonymous):